OpenStudy (anonymous):

Graph the equation by plotting 3 points -5x - 3y=-7

OpenStudy (anonymous):

-3y = 5x - 7 3y = -5x +7 now put x = 2 3y = -10 + 7 y = -3/3 = -1 first point is (2,-1) now put x = -1 3y = 5 + 7 y = 12/3 = 4 second point (-1,4)

OpenStudy (anonymous):

First rearrange the equation to the form y=mx+c. \[-5x-3y=-7\rightarrow5x+3y=7\rightarrow3y=7-5x \rightarrow y=\frac{7}{3}-\frac{5x}{3}\] Then you can read off the y intercept which is c or just plugin values for x to get 3 points

OpenStudy (anonymous):

Here is a graph so you can check your points

OpenStudy (anonymous):

now put x = 5 3y = -25 + 7 y = -18/3 = -6 so third point is (5,-6)

OpenStudy (anonymous):

@MiMi You can plot the three points given by me and join them by a line to get the graph....

OpenStudy (anonymous):

should the line match?

OpenStudy (anonymous):

okhay thanks it was right

OpenStudy (anonymous):

match what ??

OpenStudy (anonymous):

Are you ok now?

OpenStudy (anonymous):

yes thank you guys!

OpenStudy (anonymous):

Sweet :D

OpenStudy (anonymous):

harkirat can you help me again please?

OpenStudy (anonymous):

sure post yr question.... :)